# Mapping Regional Disparities in Discounted Grocery Products

### v1

@@ Results
To better understand the environmental and nutritional impact of these differences, we have mapped each available product to its corresponding Nutri-Score (72 validated products) and Environmental score (66 validated products), which respectively rate nutritional quality (from A for the healthiest to E for the least healthy) and environmental footprint (from A++ for the lowest impact to E for the highest impact). Figure 3 b shows how categories B and D are the most prevalent across all community types, indicating that the healthiest products (A) are under-represented in promotions [13]. Critically, poor nutritional scores (D+E combined) reach 0.58 in Countryside, 0.37 in Capital Region, and 0.25 in Metropolitan Areas, which presents the most balanced profile with a relatively even distribution across all score categories. In contrast, rural areas exhibit both higher promotional frequencies for high-quality foods and elevated discount rates for low-quality options, suggesting a wider but potentially riskier dietary range [6, 52]. On the other hand, Figure 3 c suggests a complementary picture regarding the Environmental score, where mid-grade (B) dominates overall, and the Countryside demonstrates environmentally positive products. However, meat products are not mapped, despite being well known to exert a disproportionately large environmental impact [18]. Notably, the Metropolitan Areas and Capital Region have a distribution shifted towards less environmentally positive products (C-D).
@@ Discussion
Human interactions with food systems generate system-wide challenges [24], creating complex feedback loops between dietary behaviors, environmental impacts, and social outcomes that reinforce one another [29, 27, 19, 20, 22]. Our analysis reveals regional differences in the on sale near-expiry products across Denmark, demonstrating that spatial context shapes both the composition and intensity of near-expiry product offerings [44]. The bipartite approach identifies three distinct retail communities – Countryside, Metropolitan Areas, and Capital Region – each exhibiting unique profiles in meat, dairy, convenience, nutritional, and environmental product assortments [63, 64]. Our findings highlight the need for food waste reduction strategies that are sensitive to local retail dynamics [39], as neglecting such spatial differences may reinforce supply chain inefficiencies and amplify environmental burdens.
@@
Overall, our results underscore the strong role of geography in shaping human behavior patterns [40, 38, 44, 48]. Food choices are grounded in local social and spatial contexts, influenced by the organization of urban and rural areas and how promotions are tailored to these settings. The regional differences in per-capita on sale frequencies and product composition patterns highlight that supply-demand mismatches vary across locations, emphasizing that effective food waste reduction strategies must be tailored to local retail dynamics rather than applied uniformly [69]. Retail patterns shift across regions and lifestyle clusters, showing how deeply place and daily decisions are linked [46]. Food waste, then, is the sum of these everyday choices – highlighting the need for strategies that tackle both geographic disparities and individual behavior to drive sustainable consumption [50].
@@ Methods
We have retrieved Data from the Salling Group Anti Food Waste API, encompassing both store and product information (See Supplementary Materials, Section 1 for all the details). Store data include geographic coordinates, unique identifiers for each location and store brand. Product data capture items on discount offers, including EAN codes, product categories, stock levels, and temporal information such as discount start and end times. Additionally, we have mapped each product to its corresponding Nutri-Score and environmental score using OpenFoodFacts, an open database of food products from around the world that aggregates ingredient lists, nutritional information, and sustainability metrics. Nutri-Score is a front-of-pack labeling system that grades foods from A (healthiest) to E (least healthy) based on their nutritional composition, while the environmental score quantifies the product’s ecological footprint across factors like greenhouse gas emissions, water usage, and land use [70]. We have collected Data during three predefined periods (2024-08-08 to 2024-10-10, 2025-03-05 to 2025-04-10, and 2025-05-01 to 2025-06-21) and analyses are limited to these observation windows. When multiple discount offers existed for the same product on the same day at a given store, only the record with the lowest price has been retained. This processing approach results in a structured dataset that captures the daily availability of discounted products across the Netto store network throughout 153 days. We have computed the shortest road‐network distance between every pair of stores using the Open Source Routing Machine (OSRM) on OpenStreetMap data [71], and obtained municipality boundaries and resident population figures for Copenhagen’s level‐2 administrative units (municipality level) from the Copernicus Global Degree of Urbanisation Classification dataset [72].
@@
To identify spatially coherent retail clusters, we have employed Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN), a density-based clustering algorithm that extends DBSCAN by converting it into a hierarchical clustering framework [58]. HDBSCAN constructs a hierarchy of clusters by varying the density parameter and extracting stable clusters across multiple density levels. We have used the pre-computed shortest path distance matrix between all store locations and set the minimum cluster size parameter to 10 stores to ensure that identified clusters represent meaningful retail agglomerations rather than isolated store pairs. The clustering procedure identified five distinct spatial clusters across Denmark’s retail landscape, grouping part of Sjælland and all of Hovedstaden into one cluster, combining a mix of Midtjylland and part of Syddanmark in another, capturing a separate part of Syddanmark in a third, forming clusters corresponding mainly to Nordjylland and Sjælland, while noise is scattered primarily across Jutland. To assess clustering quality, we have computed three complementary metrics: the Davies-Bouldin Index (0.44), which measures the average similarity ratio between clusters where lower values indicate better separation; the Silhouette Score (0.65), quantifying how well-separated clusters are with values approaching 1.0 indicating optimal clustering; and Maximum Persistence (0.364), representing the stability of clusters across the hierarchical structure where lower values suggest less robust cluster boundaries.
To quantify promotional composition differences between stores, we have computed Bray-Curtis dissimilarity indices for all store pairs based on their category-specific promotion frequencies. The Bray-Curtis index [60] is defined as
@@
where $x_{ik}$ and $x_{jk}$ represent the abundance of product category $k$ in stores $i$ and $j$, respectively, and $n$ is the total number of categories (See Supplementary Materials, Section 2 for the dendrogram). Abundance for each product category is calculated as the cumulative number of products on sale multiplied by the number of days they were offered. We have assessed the relationship between spatial proximity and product similarity using Mantel correlation analysis across multiple distance scales. The Mantel test evaluates whether two distance matrices are significantly correlated by computing Pearson correlations between corresponding elements and testing significance through permutation [61]. We have constructed a cumulative correlogram by calculating Mantel correlations at increasing distance thresholds, enabling detection of scale-dependent spatial autocorrelation patterns in promotional composition similarity. To statistically validate that identified spatial clusters correspond to genuinely distinct product composition strategies, we have performed Permutational Multivariate Analysis of Variance (PERMANOVA) using the Bray-Curtis dissimilarity matrix and the spatial clusters. PERMANOVA is a non-parametric test that evaluates whether groups differ in multivariate composition by partitioning distance matrices and calculating pseudo F-statistics. Group labels are permuted 1000 times to generate a null F-distribution, and significance is assessed by comparing the observed F to this permutation-based distribution, avoiding assumptions about data distribution.
@@
To identify products driving regional retail differentiation, we have built a bipartite network linking stores $s$ to products $p$ based on the cumulative number of days each product was on sale, $X_{s,p}$. First, we have applied RCA filtering to identify meaningful store-product associations. RCA evaluates whether a store puts on offer certain products more often than would be expected based on overall market trends [65]. It is calculated as the ratio of a store’s share of a particular product to the product’s share of total offerings across all stores:
@@
The RCA filtering yields an unweighted bipartite network that can be validated via the Bipartite Configuration Model [62, 63, 64]. BiCM defines a maximum-entropy null ensemble constrained by the observed degree sequences – that is, the number of unique products connected to each store and vice versa. The model assigns connection probabilities
@@
We perform community detection on the resulting projections using Infomap, which identifies modules by modeling information flow on the network as a random walk and optimizing a map equation [68], and Louvain, which greedily maximizes modularity [67] – a measure of the density of edges inside communities versus edges between communities. The two resulting partitions were then compared using the Adjusted Rand Index (ARI) to assess their similarity based on the contingency matrix and statistical significance via permutation testing.
Percolation on the weighted (store-store) network is performed by using travel distance as the link weight. At a given distance threshold, only edges shorter than that threshold are retained (equivalently, distances can be normalized to bond-open probabilities), allowing connected stores to form clusters that merge as the threshold increases [40].

### v2

@@ Results
To better understand the environmental and nutritional impact of these differences, we have mapped each available product to its corresponding Nutri-Score (72 validated products) and Environmental score (66 validated products), which respectively rate nutritional quality (from A for the healthiest to E for the least healthy) and environmental footprint (from A++ for the lowest impact to E for the highest impact). Figure 3 b shows how categories B and D are the most prevalent across all community types, indicating that the healthiest products (A) are under-represented in promotions [muncke_health_2025]. Critically, poor nutritional scores (D+E combined) reach 0.58 in Countryside, 0.37 in Capital Region, and 0.25 in Metropolitan Areas, which presents the most balanced profile with a relatively even distribution across all score categories. In contrast, rural areas exhibit both higher promotional frequencies for high-quality foods and elevated discount rates for low-quality options, suggesting a wider but potentially riskier dietary range [ravandi_prevalence_2025, johansen_food_2006]. On the other hand, Figure 3 c suggests a complementary picture regarding the Environmental score, where mid-grade (B) dominates overall, and the Countryside demonstrates environmentally positive products. However, meat products are not mapped, despite being well known to exert a disproportionately large environmental impact [springmann_analysis_2016]. Notably, the Metropolitan Areas and Capital Region have a distribution shifted towards less environmentally positive products (C-D).
@@ Discussion
Human interactions with food systems generate system-wide challenges [schneider_state_2023], creating complex feedback loops between dietary behaviors, environmental impacts, and social outcomes that reinforce one another [filimonau_exploratory_2017, huang_food_2021, wang_food_2025, anshassi_improving_2025, macdiarmid_food_2022]. Our analysis reveals regional differences in the near-expiry discounted products across Denmark, demonstrating that spatial context shapes both the composition and intensity of near-expiry product offerings [lengyel_role_2020]. The bipartite approach identifies three distinct retail communities – Countryside, Metropolitan Areas, and Capital Region – each exhibiting unique profiles in meat, dairy, convenience, nutritional, and environmental product assortments [saracco_inferring_2017, cimini_statistical_2019]. Our findings highlight the need for food waste reduction strategies that are sensitive to local retail dynamics [seto_hidden_2016], as neglecting such spatial differences may reinforce supply chain inefficiencies and amplify environmental burdens. While general interventions, such as the phasing out of quantity discounts, have been effective in Denmark, our results suggest that further reductions could arise from regionally differentiated policies – for example, targeting single-household packaging sizes in capital regions to reduce convenience food waste, while focusing on cold-chain inventory optimization for meat products in rural areas [halloran_addressing_2014].
@@
Overall, our results underscore the strong role of geography in shaping human behavior patterns [arcaute_cities_2016, dong_social_2018, lengyel_role_2020, sapienza_exposure_2023]. Food choices are grounded in local social and spatial contexts, influenced by the organization of urban and rural areas and how promotions are tailored to these settings. The regional differences in per-capita on sale frequencies and product composition patterns highlight that supply-demand mismatches vary across locations, emphasizing that effective food waste reduction strategies must be tailored to local retail dynamics rather than applied uniformly [dogan_cities_2023]. Retail patterns shift across regions and lifestyle clusters, showing how deeply place and daily decisions are linked [pan_urban_2013]. Food waste, then, is the sum of these everyday choices – highlighting the need for strategies that tackle both geographic disparities and individual behavior to drive sustainable consumption [otto_social_2020].
@@ Methods
We have retrieved Data from the Salling Group Anti Food Waste API, encompassing both store and product information (See Supplementary Materials, Section 1 for all the details). Store data include geographic coordinates, unique identifiers for each location and store brand. Product data capture items on discount offers, including EAN codes, product categories, stock levels, and temporal information such as discount start and end times. Additionally, we have mapped each product to its corresponding Nutri-Score and environmental score using OpenFoodFacts, an open database of food products from around the world that aggregates ingredient lists, nutritional information, and sustainability metrics. Nutri-Score is a front-of-pack labeling system that grades foods from A (healthiest) to E (least healthy) based on their nutritional composition, while the environmental score quantifies the product’s ecological footprint across factors like greenhouse gas emissions, water usage, and land use [andreani_nutri-score_2025]. We have collected Data during three predefined periods (2024-08-08 to 2024-10-10, 2025-03-05 to 2025-04-10, and 2025-05-01 to 2025-06-21) and analyses are limited to these observation windows. When multiple discount offers existed for the same product on the same day at a given store, only the record with the lowest price has been retained. This processing approach results in a structured dataset that captures the daily availability of discounted products across the Netto store network throughout 153 days. We have computed the shortest road‐network distance between every pair of stores using the Open Source Routing Machine (OSRM) on OpenStreetMap data [huber_osrmtime_2016], and obtained municipality boundaries and resident population figures for Copenhagen’s level‐2 administrative units (municipality level) from the Copernicus Global Degree of Urbanisation Classification dataset [european_commission_joint_research_centre_ghsl_2023].
@@
To identify spatially coherent retail clusters, we have employed Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN), a density-based clustering algorithm that extends DBSCAN by converting it into a hierarchical clustering framework [hutchison_density-based_2013]. HDBSCAN constructs a hierarchy of clusters by varying the density parameter and extracting stable clusters across multiple density levels. We have used the pre-computed shortest path distance matrix between all store locations and set the minimum cluster size parameter to 10 stores to ensure that identified clusters represent meaningful retail agglomerations rather than isolated store pairs. The clustering procedure identified five distinct spatial clusters across Denmark’s retail landscape, grouping part of Sjælland and all of Hovedstaden into one cluster, combining a mix of Midtjylland and part of Syddanmark in another, capturing a separate part of Syddanmark in a third, forming clusters corresponding mainly to Nordjylland and Sjælland, while noise is scattered primarily across Jutland. To assess clustering quality, we have computed three complementary metrics: the Davies-Bouldin Index (0.44), which measures the average similarity ratio between clusters where lower values indicate better separation; the Silhouette Score (0.65), quantifying how well-separated clusters are with values approaching 1.0 indicating optimal clustering; and Maximum Persistence (0.364), representing the stability of clusters across the hierarchical structure where lower values suggest less robust cluster boundaries.
To quantify promotional composition differences between stores, we have computed Bray-Curtis dissimilarity indices for all store pairs based on their category-specific promotion frequencies. The Bray-Curtis index [bray_ordination_1957] is defined as
@@
where $x_{ik}$ and $x_{jk}$ represent the abundance of product category $k$ in stores $i$ and $j$, respectively, and $n$ is the total number of categories (See Supplementary Materials, Section 2 for the dendrogram). Abundance for each product category is calculated as the cumulative number of products on sale multiplied by the number of days they were offered. We have assessed the relationship between spatial proximity and product similarity using Mantel correlation analysis across multiple distance scales. The Mantel test evaluates whether two distance matrices are significantly correlated by computing Pearson correlations between corresponding elements and testing significance through permutation [diniz-filho_mantel_2013]. We have constructed a cumulative correlogram by calculating Mantel correlations at increasing distance thresholds, enabling detection of scale-dependent spatial autocorrelation patterns in promotional composition similarity. To statistically validate that identified spatial clusters correspond to genuinely distinct product composition strategies, we have performed Permutational Multivariate Analysis of Variance (PERMANOVA) using the Bray-Curtis dissimilarity matrix and the spatial clusters. PERMANOVA is a non-parametric test that evaluates whether groups differ in multivariate composition by partitioning distance matrices and calculating pseudo F-statistics. Group labels are permuted 1000 times to generate a null F-distribution, and significance is assessed by comparing the observed F to this permutation-based distribution, avoiding assumptions about data distribution.
@@
To identify products driving regional retail differentiation, we have built a bipartite network linking stores $s$ to products $p$ based on the cumulative number of days each product was on sale, $X_{s,p}$. First, we have applied RCA filtering to identify meaningful store-product associations. RCA evaluates whether a store puts on offer certain products more often than would be expected based on overall market trends [balassa_revealed_1989]. It is calculated as the ratio of a store’s share of a particular product to the product’s share of total offerings across all stores:
@@
The RCA filtering yields an unweighted bipartite network that can be validated via the Bipartite Configuration Model [saracco_randomizing_2015, saracco_inferring_2017, cimini_statistical_2019]. BiCM defines a maximum-entropy null ensemble constrained by the observed degree sequences – that is, the number of unique products connected to each store and vice versa. The model assigns connection probabilities
@@
We perform community detection on the resulting projections using Infomap, which identifies modules by modeling information flow on the network as a random walk and optimizing a map equation [rosvall_maps_2008], and Louvain, which greedily maximizes modularity [blondel_fast_2008] – a measure of the density of edges inside communities versus edges between communities. The two resulting partitions were then compared using the Adjusted Rand Index (ARI) to assess their similarity based on the contingency matrix and statistical significance via permutation testing.
Percolation on the weighted (store-store) network is performed by using travel distance as the link weight. At a given distance threshold, only edges shorter than that threshold are retained (equivalently, distances can be normalized to bond-open probabilities), allowing connected stores to form clusters that merge as the threshold increases [arcaute_cities_2016].